Problem 16-51 (Algo) Solve for Master Budget Given Actual Results (LO 16-2, 4) A new accounting intern at Gibson Corporation lost the only copy of this period's master budget. The CFO wants to evaluate performance for this period but needs the master budg

A new accounting intern at Gibson Corporation lost the only copy of this period's master budget. The CFO wants to evaluate performance for this period but needs the master budget to do so. Actual results for the period follow.
  





Sales volume
150,000units
Sales revenue$1,008,000
Variable costs


Manufacturing
221,760
Marketing and administrative
90,720
Contribution margin$695,520
Fixed costs


Manufacturing
277,000
Marketing and administrative
153,700
Operating profit$264,820

  

The company planned to produce and sell 120,000 units for $6.00 each. At that volume, the contribution margin would have been $504,000. Variable marketing and administrative costs are budgeted at 10 percent of sales revenue. Manufacturing fixed costs are estimated at $2.40 per unit at the normal volume of 120,000 units. Management notes, "We budget an operating profit of $1.00 per unit at the normal volume."

  

Required:

a. Construct the master budget for the period.
b. Prepare a profit variance analysis.
Answer #1

A)

sales volume   120,000

sales revenue= 120,000* 6= 720,000

variable costs:

manufacturing  720,000- 504,000- 72,000= 144,000

Marketing and administrative 720,000*10%= 72,000

Contribution margin  720,000-144,000-72,000= 504,000

Fixed costs:    

Manufacturing    120,000* 2.40= 288,000

Marketing and administrative 504,000- 288,000- 120,000= 96,000

operating profit 504,000- 288,000- 96,000= 120,000


B)

flexible budget column

sales revenue 150,000* 6= 900,000

market and admin= 900,000* 10%= 90,000

manufacturing 144,000/120,000* 150,000= 180,000

contribution margin= 900,000- 180,000- 90,000= 630,000

fixed cost: manufacturing 288,000

market and admin  96,000

operating profit  630,000- 288,000- 960,000= 246,000


Master budget column same as in part A


sales activity variance column

subtract flexible budget from master budget except from fixed cost section


manufacturing variances column

variable cost

manufacturing= 221,760(actual column)- 180,000(flexible)= 41,760 U

CM= 41,760 U

fixed cost: manufacturing 277,000- 288,000= 11,000 F

operating profit 41,760-11,000= 30,760 U


market and admin variances column

variable cost market +admin= 90,720(actual)- 90,000(flexible)= 720 U

CM= 720 U

fixed cost market+admin= 153,700- 96,000= 57,700 U

operating profit= 720+ 57,700= 58,420 U


sales price variance column

sales revenue 1,008,000(actual)- 900,000(flexible)= 108,000 F

CM= 108,000 F

operating profit= 108,000 F